Grade 12th passPhysical Chemistry

2 liquid x and y form of a solution containing 1 mole of x and 3 mole of y is 550nm of Hg if one mole of y is further added to this solution incrrased by 10 nm of Hg calculate the vapour pressure of x and y in pure stable in pure state?

ankit singh
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear Ankit
 
P = x*Pa + y* Pb
(1/1+3)Pa  + (3/1+3) Pb = 550
Pa + 3 Pb = 2200.      .....(i)
(1/1+4) Pa + (4/1+4) Pb = 560
Pa + 4Pb = 2800.         ......(ii)
 
Solve (i) and (ii)
Pa = 400 mm Hg
Pb = 600 mm Hg
 
 
Regards
Arun (askIITians forum expert)